Home‎ > ‎Reports‎ > ‎Crew100B‎ > ‎

day06


Crew 100B Log Book for March 3 2011

Photos of the day





Commander's Check-In Report

Balwant Rai reporting

Experimental Status: All the medical, habitability, astrobiological and food experiments began today with unique results.

Time Departed/Returned from EVA: three EVAs performed

Time Departed/Returned from EVA: As per EVA reports

The meals today were just satisfying since it was a cooking day.

EVA’s: We completed three successful EVA’s.

Number Of reports: 2 science reports, 1 engineering report, 1 journalist report, 3 EVA reports and 1 chef report.

Plan for tomorrow: will initiate the medical experiment an habitability again on all crew members. Interpretation of the Sun by means of the radio telescope will also commence. The study of the rover’s maneuverability as well as the potential use of the drone, weather pending, will take place. Soil samples will yet again be collected during three EVA’s tomorrow. One will be on foot, and two will use the ATV’s. Tomorrow will be a non-cooking day.


Commander's Report

Balwant Rai Reporting

Today we commenced the day with our continuing medical experiments and our habitability experiment.

The crew’s medical health will continue to be monitored by our health and safety officer for the remainder of our stay here on Mars. Breakfast was very good and delicious. Our crew had their saliva samples taken (via Versi Sal 1 from Oasis Diagnostics), vital parameters monitored (via the Zephyr Bioharness), and their heart rate variability measured (via NERV Express 4.2 software). I also performed the cognitive study (via CogState Research software) on all crew members, as well.

Our crew scientists continued to work through problems with starting observations of the sun via the radio telescope. Three EVA’s were performed to sites of interest. Today is a cooking day, so I am pretty sure our chief and associate chef is busy finding something that might taste very delicious. Tomorrow, we will continue our medical, habitability and astrobiological experiments, as well as continue getting the rover working well, and potentially getting the radio telescope up and working.




Engineering Report

Matthieu Ansart Reporting

Generator/Electricity:

· Kitty (Diesel Generator) (Used? Yes/No): Yes

· Honey (Gas Generator) (Used? Yes/No): No

· Battery Charge Level: Checked by DG

o 10.59 V everstart

o 12.37 V armor plate

o 13.32 V green

Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):

1 generator run, all is good


Fuel Status:

· Diesel (Full/0.75/0.5/0.25/Empty): between 0.75 and 0.5

· Propane Tank (Full/0.75/0.5/0.25/Empty): 0.38

· Gasoline Tank (Full/0.75/0.5/0.25/Empty): 2 containers full

Notes/Comments:


Oil:

· Kitty Oil Quantity (# of Quarts in storage): Leave it to DG

· Honey Oil Quantity (# of Quarts in storage): Leave it to DG

· ATV Oil Quantity (# of Quarts in storage): 6 quarts on the engineering work bench

· Car Oil Quantity (# of Quarts in storage):


Water Status:

· Outside Potable Water Tank Level (inches from bottom): 15

· Trailer Potable Water Tank Level (inches from bottom): 1/3 Full

· Hab Potable Water Tank Level (inches from bottom): 13

· Potable Water Meter Reading


Notes/Comments: Yesterday in the evening, there was a leak outside above the outside door, after switch on the water heater. The leak stopped some minutes after I switched off the water heater.



GreenHab:

· Condition of plants in Tank 1: green water

· Condition of plants in Tank 2: green plants and snails

· Condition of plants in Tank 3: green plants

· Greenhab notes/ comments: Aquatic Tank 1 is full. Aquatic Tank 2 has begun to be filled. In Aquatic Tank 2, we find some sort of snails with duckweeds. Is it normal? Do we have something to do with them?


Observatory:

· Telescope Used? (Yes/no) No

Observatory Notes


Transportation:

Hab Car (NOT RENTAL)

· Used (yes/no): No

· Oil Checked (yes/no): No

· Tire Status: Not checked

· Odometer at end of day: 123973

· Notes/Comments on Hab Car:


ATV 1 (Opportunity):

· Used: Yes

· Oil Checked: no

· Fuel Consumed:Replaced

· Tire Status: good

· Comments:


ATV 2 (Spirit):

· Used: No

· Oil Checked: No

· Fuel Consumed:

· Tire Status: the two left are deflated

· Comments: we checked this ATV. The wheel axis on the back left seems to be twisted.


ATV 3 (Viking I):

· Used: Yes

· Oil Checked: no

· Fuel Consumed: Replaced

· Tire Status: good

· Comments:


ATV 4 (Viking II):

· Used: Yes

· Oil Checked: no

· Fuel Consumed:Replaced

· Tire Status: good

· Comments:


Heating and Ventilation:

· Thermostat setting upstairs (Farenheit): 65

· Actual temperature upstairs (Farenheit): 72

· Thermostat setting downstairs (Farenheit): 55

· Actual temperature downstairs (Farenheit): 63


Computers/Networking Infrastructure

Notes/comments on internet/computers?:


General Engineering

1. Notes/Comments:


Science Report (Medical)

Jasdeep Kaur Reporting

Time: 15:00-Blood pressure and pulse rate monitored before and after physical work.

Crew Members: Science Team: 3 crew members Quentin Bourges,Matthieu Ansart and Jasdeep Kaur

Location: Hab

Objective: To monitor the blood pressure, pulse rate and heart rate after physical work to find if there exist some major changes.

Observation: Three crew members were monitored for their blood pressures. The blood pressures and heart rate was also monitored by Zephyr technology. Pulse was checked manually and electronically prior and after the physical work.

Results: Blood Pressure, pulse rate and heart rate was not significantly affected after physical work. The physical work was not very strenuous. The physical work included execution of some experiments of crew members and also EVA activities.

Lessons learned: Not recommended to wear full sleeve while giving blood pressure findings as it could give some errors..


Science Report (Bio-Medical)

Balwant Rai reporting

Heart rate and arterial pressure in EVA Participants of Mars Analogue Mission

Spaceflight causes adaptive changes in cardiovascular physiology such as post-flight orthostatic intolerance, that can have deleterious effects on astronauts. In-flight cardiovascular data are difficult to obtain, and results have been inconsistent. A journey to Mars will take months to complete and as such an appropriate exercise regime must be in place to ensure the Safety and productivity of the crew. To the best my knowledge no evaluation made on effect of simulated environment on heart rate and arterial pressure involved in Mars analogue EVA activity. Hence this study is planned to study the effects of EVA on heart rate and arterial pressure . All crews’ members were selected for this study. Heart Rate and arterial pressure were measured before and after EVA by Zephyr BioHarness. Heart rate, diastolic pressure, variability of heart rate and diastolic pressure levels were significantly increased. So astronauts should be required to participate in physical activity to minimize these adverse effects.


EVA 16 Report

Jasdeep Kaur Reporting

EVA # 16, 10:55-13:00

Crew Members: Matthieu Ansart, Quentin Bourgesr and Jasdeep Kaur

Transit Mode: ATVs

Objectives: Geo-Mission: to locate dinosaur's dig to find some bones and fossils

Results: At 9 am three crew members departed for EVA on ATV. This was a planned ATV and crew members carried all required equipments to locate fossils and if possible bones. Crew members went in a huge dig and located big feet of cattle in dig. Also, it gave a clear idea of water flow or river in past, which is totally dried at present. Also, we could find the area with really amazing stuff. special features in the area was the big holes which can be niche for any kind of reptiles, so crew members had to really take care of not pressing hard on such holes and not to disturb by tapping over those. It was really adventurous to walk in middle of totally dried river. We also climbed the high mountains over to locate some findings. we also found the stones and the high areas appeared almost hollow from inside, so this was another sign of some species living there which could be dangerous. The stones at some places almost looked similar to texture of bone, but of course it was due to water flow in past and these were good evidence of sediments. We could find some chameleons running around.

Lessons Learned: big holes definitely could have some dangerous reptiles, so, its not recommended to visit that area in late evening and also not to push hard walking walking and to be very cautious of everything around.


EVA 17 Report

Matthieu Ansart Reporting

EVA # 17, 12:00-12:15

Crew Members: Matthieu Ansart, Quentin Bourges

Transit Mode : Foot

Objectives : Try to understand what is the problem with the ATV Spirit and try to fix it.

Results : The two left tires are deflated. However the main problem is concerning the back axis. The back left wheel doesn't turn properly. The axis of rotation is twisted. So we should have a spare one to fix the ATV until that moment, Spirit is definitely useless.


EVA 18 Report

Crystal Latham Reporting

EVA # 18, 16:00-16:45

Crew Members: Rachel Dompnier, Crystal Latham, Balwant Rai

Site Location: Confidential

Transit Mode: ATVs

Objectives: Geo-Mission: to locate dinosaur's dig to find some bones and fossils

Results: We almost made it to the site, but it began to rain. So we had to turn around and go back to the HAB.

Lessons Learned: To check the weather report prior to going on long EVA’s. Having a weather station here at MDRS so weather info for the area can be known by the crew more accurately.


Chef's Report

Quentin Bourges Reporting

1a) Today was

- a cooking day

1b) Was today a special day celebrated at the main meal with special food or activity? If so, what was special about today and what food(s) were prepared to mark the celebration?

No

2) List the foods served at today's main meal (usually the evening meal), giving full names of each.

Leftovers from yesterday and

Onions, green peas, ground cinnamon, garlic salt, ground ginger, ground turmeric, chili and spinach.

3) List any main meal foods not finished by the crew (leftovers stored for later use or discarded)

It was so good, there wasn’t any leftover.

4) Recipes for anything more complicated than following package directions exactly, or rehydrating.

Cook some brown rice, add dry milk with hot water and sugar. Eat it while it’s still warm!

5) Lessons learned (if any)

No

6) Comments/questions for Kim and Jean, the food study investigators

No


Journalist Report

Rachel Dompnier reporting

Crew 100B’s members had a very enjoyable breakfast this morning. Some went on an EVA whereas others stayed at the HAB to perform medical experiments.

Today was too windy to go out with the drone. The weather was not cooperating and we had to abort an EVA due to some unpredictable rain.

A few of the crewmembers tried to fix the broken ATV, but they were unsuccessful.

We also saw a local Martian today; his name was DJ and was very friendly.

Today was a cooking day where we tried to cook brown rice but it was a total failure. Later, when we started to cook dinner, we attempted to make bread but luck seems not to be with us today…